OpenStudy (anonymous):

How would you simplify e^x+ln3 ?

OpenStudy (anonymous):

Guide me please, no direct answers

OpenStudy (anonymous):

e^(x+ln3)

OpenStudy (australopithecus):

Sorry this is easy

OpenStudy (australopithecus):

remember the rule, A^(b + c) = A^bA^c

OpenStudy (anonymous):

The answer is 3e^x Wondering how this was derived.

OpenStudy (australopithecus):

Also remember the rule e^ln(x) = x ln(e) = 1

OpenStudy (australopithecus):

One more rule you need to remember ln(x^a) = aln(x)

OpenStudy (australopithecus):

Can you solve it now?

OpenStudy (australopithecus):

Sorry I'm pretty exhausted can't believe I spaced on this

OpenStudy (jdoe0001):

hehe

OpenStudy (anonymous):

Not really.. The x is confusing me

OpenStudy (jdoe0001):

\(\bf \Large{ e^{x+ln(3)}\implies e^{x+log_e(3)}\implies e^x\cdot e^{log_e(3)}\\ \quad \\ \textit{log cancellation rule of } {\color{red}{ a}}^{log_{\color{red}{ a}}x}=x\quad thus\\ \quad \\ e^x\cdot {\color{red}{ e}}^{log_{\color{red}{ e}}(3)}\implies e^x\cdot \square ?}\)

OpenStudy (anonymous):

Still don't understand it. Will need to watch a video again. I hate logs.

OpenStudy (australopithecus):

Logs are easy, just look at the rules I posted

OpenStudy (jdoe0001):

heeh lemme rewrite the rule withuot the "x"

OpenStudy (australopithecus):

Dont get frustrated you can do this :)

OpenStudy (jdoe0001):

\(\bf \Large{ e^{x+ln(3)}\implies e^{x+log_e(3)}\implies e^x\cdot e^{log_e(3)}\\ \quad \\ \textit{log cancellation rule of } {\color{red}{ a}}^{log_{\color{red}{ a}}\square }=\square \quad thus\\ \quad \\ e^x\cdot {\color{red}{ e}}^{log_{\color{red}{ e}}(3)}\implies e^x\cdot \square ?}\)

OpenStudy (anonymous):

Ok so the ln^3 becomes = 3 And we are left with the e^x So 3e^x Is that all?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

Ok cool thanks you guys will need some more practice!

OpenStudy (jdoe0001):

yw

OpenStudy (mathmale):

@intraining: Part of the difficulty with this problem stems from the ambiguity in e^x + ln 3. If one were to take this expression literally, and graph the result, the graph would look lik3e that of y=e^x shifted upward by the constant quantity ln 3. On second thought, it occurred to me that you meant e^(x+ln 3). This is an entirely different problem. Since \[e ^{a+b}=e ^{a}e ^{b}\] \[e ^{x+\ln 3}=e ^{x}e ^{\ln 3}=3e ^{x}\] When in doubt, please use parentheses!

OpenStudy (anonymous):

Thank you mathmale!

OpenStudy (mathmale):

My great pleasure!!